(1 pt) Because of the relatively high interest rates, most consumers attempt to pay off their credit card bills promptly. However, this is not always possible. An analysis of the amount of interest paid monthly by a banks Visa cardholders reveals that the amount is normally distributed with a mean of 27 dollars and a standard deviation of 6 dollars. A. What proportion of the banks Visa cardholders pay more than 29 dollars in interest? Proportion 0.3707 B. What proportion or the banks Visa cardholders pay more than 38 dollars in interest? Proportion0.0334 C. What proportion of the banks Visa cardholders pay less than 16 dollars in interest? Proportion0.0334 D. What interest payment is exceeded by only 21% of the banks Visa cardholders? Interest Payment31.836

Please do part A
Answer #1

mu= 27

sigma= 6

X= 29

Z=(X-mu)/sigma

=(29-27)/6

0.333333

probability =P(Z>0.3333)

=0.369441

rounded =0.3694

........................

if you have doubt ask in comment give thumbs up if you like work